OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2))

hero (hero):

f(g(x)) = g(x) + 8 = (x^2 - 6x - 7) + 8 = x^2 - 6x - 7 + 8 = x^2 - 6x + 1 f(g(2)) = (2)^2 - 6(2) + 1

OpenStudy (anonymous):

Let f(x) = 2x + 2. Solve f-1(x) when x = 4.

hero (hero):

Find the inverse of f(x). Let y = f(x) y = 2x + 2 x = 2y + 2 x - 2 = 2y \(\dfrac{x - 2}{2} = y\) \(\dfrac{x - 2}{2} = f^{-1}(x)\)

hero (hero):

Finish it

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1))

OpenStudy (anonymous):

Plug -1 in for x in the g(x) equation. then plug that answer in for x in the f(x) equation

OpenStudy (anonymous):

i got it, it was -3